1. Q = a + bP

a. How is “a” called? Linear demand


b. What is the whole equation representing usually in the context of the theory of consumption?


c. How much is the change in Q per unit change in P?

d. For Q to be equal to “a,” what must be the value of P?

e. What is the expected shape of the curve formed by the points of Q and P in a graph?

2. Consider the two equations (10 points):

Q = a + bP

Q = c – dP

Derive the values of Q and P.



1
Expert's answer
2021-04-29T10:21:26-0400

1. a. "a" is the quantity demanded when the price P = 0.

b. The whole equation representing usually in the context of the theory of consumptionis called the demand function.

c. The change in Q per unit change in P is b.

d. For Q to be equal to “a,” the value of P must be:

a = a + bP,

bP = 0,

P = 0.

e. The expected shape of the curve formed by the points of Q and P in a graph is a line.


2. If Q = a + bP and Q = c – dP, then:

a + bP = c – dP,

P(b + d) = c - a,

"P = (c - a)\/(b + d)."

"Q = a + b\u00d7(c - a)\/(b + d)."